OpenStudy (anonymous):

How would you identify the empirical formula of a compound CuCla bH2O. You have at your disposal, an electronic balance and any other equipment you may need.

OpenStudy (anonymous):

I need help on finding a.

OpenStudy (aayushi.somani):

Mass of water = 1.0 - 0.76 = 0.24 g = 0.013 moles, OK. 0.43 g Cu x (1 mole Cu / 63.5 g Cu) = 0.0068 moles Cu 0.33 g Cl x (1 mole Cl / 35.5 g Cl) = 0.0093 moles Cl. Dividing by the smallest (0.0068), we get a Cu / Cl ratio of 1/ 1.36 which is 3:4 so your Cu3Cl4 isn't unreasonable (but I've never heard of Cu3Cl4). However, I think that it's MUCH more likely that the ratio is 1 / 1.5 and the correct formula is CuCl2. That compound has a molar mass of 134.4. 0.76 g CuCl2 x (1 mole CuCl2 / 134.4 g CuCl2) = 0.0057 moles CuCl2 0.013 moles H2O / 0.0057 moles CuCl2 = 2.3 moles H2O / mole CuCl2. The likely formula is CuCl2 x 2H2O

OpenStudy (anonymous):

You're supposed to conduct an experiment.

OpenStudy (anonymous):

How would you know all of this?

OpenStudy (aayushi.somani):

by titration. this is an experiment ot find the mass% and hence the empirical formula..!

OpenStudy (anonymous):

This is what I think please correct me if I am wrong: Method: 1. Evaporate water off from the compound by heating it in a ceramic flask, over a bunsen burner and a tri-pod. 2. Weigh the anhydrous sample of CuCla. 3. React the CuCla whith K(s) wait until no further precipitate is formed. So CuCla + K -> KCl + Cu 4. Weigh the sample of Cu. 5. Find the moles of Cu by deviding by molar mass. Subtract mass of Cu from CuCla to get mass Cla, devide by its molar mass to get moles. 6 Ratio moles Cu with moles Cl.

OpenStudy (aayushi.somani):

@AnImEfReaK is this what u think is titration????

OpenStudy (anonymous):

No I'm using metal displacement.

OpenStudy (aayushi.somani):

ohk well if u use metal displacement method then how would u find the emperical formula or the mass% ???

OpenStudy (anonymous):

mass%?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

is my method right?

OpenStudy (aayushi.somani):

mass% is the % of that element in the whole compound..!

OpenStudy (anonymous):

Ok i get that

OpenStudy (aayushi.somani):

i dont know how to find Emp. formula by metal displacemet.. but in titration i can help u!

OpenStudy (anonymous):

I would ratio the moles of Cu and Cla Cu:Cla to get the mole ratios.

OpenStudy (anonymous):

Okay can you explain step by step please?

OpenStudy (aayushi.somani):

yeah shure m sending u a link of a video in which titration is performed..!! http://www.youtube.com/watch?v=sFpFCPTDv2w

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

The video was helpful but lets say you added 50mL of CuCla and 19mL of HCL at 2 mol/L reacted, even though how would you find moles of CuCLa?

OpenStudy (aayushi.somani):

ur talking about titration?

OpenStudy (anonymous):

Yes

OpenStudy (aayushi.somani):

by molarity..!! @AnImEfReaK

OpenStudy (anonymous):

?

OpenStudy (aayushi.somani):

molarity =moles / volume..

OpenStudy (anonymous):

mol HCL = 38 mol CuCl = 50 mL * what concentration?

OpenStudy (aayushi.somani):

mol is never meseared in ml..!!

OpenStudy (anonymous):

It is n = C times V

OpenStudy (aayushi.somani):

The molecular weight of a compound is simply the sum of all the individual components. In CuCl2 x 2H2O, there are simply two additional water molecules hydrating your CuCl2. Therefore, MW (CuCl2 x 2H2O) = Cu + 2(Cl) + 2[2(H) + O] = (63.54 g/mole) + 2(35.45 g/mole) + 2[2(1 g/mole) + 16 g/mole)] = 170.44 grams/mole If you have 0.502 g of this substance, then you will have: 0.502 g/(170.44 g/mole) = 0.00295 moles of CuCl2 x 2H2O

OpenStudy (aayushi.somani):

then u have the volume and now moles as well soo now can u find the molarity??

OpenStudy (anonymous):

but how do you even know a is 2 and b is 2

OpenStudy (aayushi.somani):

i found it earlier..!!! scroll it..!!

OpenStudy (anonymous):

Cu can take on +1 2 or 3 or 4 it has many oxidation states you can't assume it takes on that form.

OpenStudy (anonymous):

I'm very confused...

OpenStudy (aayushi.somani):

there is a method of finding the empirical formula.. and by that i have found it..!! scroll these comments and u'll see the complete method i had done

OpenStudy (anonymous):

Mass of water = 1.0 - 0.76 = 0.24 g = 0.013 moles, OK. 0.43 g Cu x (1 mole Cu / 63.5 g Cu) = 0.0068 moles Cu 0.33 g Cl x (1 mole Cl / 35.5 g Cl) = 0.0093 moles Cl.

OpenStudy (anonymous):

I don't get this

OpenStudy (anonymous):

Did you separate the compound?

OpenStudy (aayushi.somani):

yeah..!! i have foound moles of each seperately

OpenStudy (anonymous):

But how? What did you do?

OpenStudy (anonymous):

You have to write steps down so not figures.

OpenStudy (anonymous):

I'll understand it better that way.

OpenStudy (aayushi.somani):

i am posting a generalised formula \[\frac{ Vacid*Macid }{Vbase*Mbase }=\frac{ no. of moles of acid \in the balanced equation }{ no. of moles of base \in the balanced equation }\]